Nitrogen gas is produced in a reaction. If the temperature of the gas is 289.5 K, the pressure at 11.7 MPa, and the volume is 12.8 L, determine the mass of the nitrogen gas, in kilogram.



1
Expert's answer
2022-04-08T11:06:03-0400

"PV=nRT"

R=8.31mol-1K-1

"n=\frac{PV}{RT}=\frac{11.7x10^6x0.0128}{8.31x289.5}=62.251 mol"

M(N2)=14x2=28 g/mol

m(N2)=28x62.251=1743.027 g=1.743kg